PART 1 Palm Corporation had total payroll of$600,000. For this payroll period, the company withheld 20% for employee income taxes. The FICA tax rate is 7.65% for wages on each employee up to their first $1 00,000 of earnings, and 1 .45% for earnings in excess of $ 1 00,000. The employer must match the employees FICA withholding with an equal amount as employer payroll taxes. In addition, the employer must pay state and federal unemployment taxes on the first $7,000 of employee wages earned. The state unemployment tax rate is 1.5% and the federal unemployment tax rate is 0.8%. Ofthe wages earned, $50,000 represented amounts earned byemployees who have already exceeded the S100,000 earnings limit, and $425,000 represented amounts earned by employees who have already exceeded the $7,000 earnings lim 1 t. REQUIRED: (1 Prepare the general journal entry, in proper form, to record the employee payroll for the period. Prepare the general journal entry, in proper form, to record the employer payroll taxes for the period. (2)
PART 2 Palm Corporation has an agreement with its sales manager that he is to receive a bonus of 7.5% of net income after the bonus and income taxes. The corporations income before deducting the bonus and income taxes is $350,000. Income taxes are 40% and the bonus is deductible for income tax purposes REQUIRED Calculate the amount of the bonus that will be paid to the sales manager. Round your answer to the nearest whole dollar and show all supporting computations Calculate the amount of the income taxes that the company will pay. Round your answer to the nearest whole dollar and show a supporting computations (2)
Answer #1

PART 1

1. General journal entry to record the employee payroll for the period
Salaries & Wages $600,000
     Withheld Income Taxes payable $120,000
     FICA tax payable $42,800
     Net Payroll payable $437,200
Working note:
FICA Tax Payable = $42075 + $725 = $42800
            = 50000 x 1.45% = $725
            = 550000 x 7.65% = $42075
2. General journal entry to record employer payroll taxes for the period
FICA Tax expenses $42,800
State unemployment tax expenses $2,625
Federal unemployment tax expenses $1,400
       FICA Tax payable $42,800
       State unemployment tax payable $2,625
       Federal unemployment tax payable $1,400
Working note:
State unemployment tax = (600000-425000) x 1.5%
                                                    = 175000 *1.5%
                                                    = 2625
Federal unemployment tax = (600000-425000) x 0.8%
                                                    = 175000 *0.8%
                                                    = 1400
PART 2
1 . Amount of Bonus to be paid to sales manager
Let B = Bonus ; T = Income tax
      B = 7.5% (350,000 - B - T)
      B = 26250 -0.075B -0.075T
Let T = Income Tax
          = 40%(350,000 - B)
          = 140000 - 0.40B
Now Substitute the equation for T in the equation for B
        B = 26250 - 0.075B - 0.075(140000 - 0.40B)
        B = 26250 - 0.075B - 10500 + 0.03B
      1.045B = 15750
       B = 15072
Proof:
Income before bonus and income tax $350,000
Less: Bonus $15,072
Net income Before income tax $334,928
Less: Income tax (40%) $133,971
Net income after bonus and taxes $200,957
Bonus as computed above
Net income after bonus and taxes $200,957
Multiply by Bonus rate 7.50%
Bonus $15,072
2. Amount of Income tax that the company will pay
    T = 40%(350000-15072)
    T = 40%(334928)
    T = $133971